14. How many multiples of 4 lie between 10 and 250?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The first multiple of 4 that is greater than 10 is 12. Next multiple will be 16. Therefore, the series formed as; 12, 16, 20, 24, … All these are divisible by 4 and thus, all these are terms of an A.P. with first term as 12 and common difference as 4. When we divide 250 by 4, the remainder will be 2. Therefore, 250 − 2 = 248 is divisible by 4. The series is as follows, now; 12, 16, 20, 24, …, 248 Let 248 be the nth term of this A.P. first term, a = 12 common difference, d = 4 an = 248 As we know, an = a+(n−1)d 248 = 12+(n-1)×4 236/4 = n-1 59 = n-1 n = 60 Therefore, there are 60 multiples of 4 between 10 and 250.